profit-maximizing firm in a competitive market is currently producing 110 units of output. It has average revenue of Rs.1000, average total cost of Rs.800, and fixed cost of Rs.20,000.

  1. What is its profit? 
  2. What is its marginal cost? 
  3. Is the efficient scale of the firm more than, less than, or exactly 100 units? 
1
Expert's answer
2021-06-07T13:14:23-0400

1.

"Profit=TR-TC\\=110(1000-800)\\=22,000"

Therefore the profit is R22,000


2.

For profit maximization firms,

"MC=MR\\"

and "MR=AR" in a competitive market

"\therefore MC=R1,000"


3.

Efficient scale of production occurs when AC=MC

here, "MC=R1,000>AC=R800"

So, the efficient scale of production of the firm is less than 110 units